what is the freezing point of a 1,77m solution of cbr4 in c6h6? given that kf is 4.90 for c6h6 the normal freezing point of c6h6 is 5.51c

Grade:12th pass

2 Answers

Sahil Yadav
2 years ago

The osmotic pressure of same solution is 0.474 atm


ΔTf=5.55.4=0.1oC
 
π=CMST(C+M=molarity)
 
ΔTf=Cm×Kf(Cm=molality)
for dilute solutions CM=Cm
 
π=KfΔTf×ST=4.90.1×0.0821×283=0.474atm
Sahil Yadav
2 years ago
you can chech this answer, i have done this same type numerical.
again if u have any doubt u can ask again i will solve the same question u have asked in doubt.